The sum of your parents age is 74. Your father is 6 years older than your mother. Write a system of equations to find the ages of both parents.
Answer by fractalier(6550)

Call their ages NOW, f and m.
Thus we have
f + m = 74 and
f = m + 6
To solve, substitute the second into the first and get
m + 6 + m = 74
2m + 6 = 74
2m = 68
m = 34
Then f = 34 + 6 = 40